OpenStudy (anonymous):

f(x) = -7/2x When f(x)=14

OpenStudy (anonymous):

Just substitute x for 14. f(x) = -7 / 2x F(14) = -7/2(14) now keep going and solve

OpenStudy (anonymous):

Which is -49 ?

OpenStudy (anonymous):

Correct.. Sorry I'm taking a test while I'm helping other too. LOL

OpenStudy (cornzyblack):

f(x) = -7/2x Substitute for x if its 14 f(14) = -7/ (2 * 14) = -7/(28) = -1/4